# Semantic-guided Search for Efficient Program Repair with Large Language Models

### v1

@@ Abstract.
Fixing software bugs is crucial yet demands significant resources from developers. Automated Program Repair (APR) is a promising solution to address this challenging task. The emergence of Large Language Models (LLMs) has opened a new era of LLM-based APR, substantially progressed the APR field further. However, conventional LLM-based Program Repair methods face significant challenges regarding memory inefficiency, hindering their scalability and effectiveness. This is largely due to the beam search utilized in the patch generation phase of LLM-based Program Repair, which often requires large beam sizes to search for more potentially good repair candidates.
In this paper, we first show that increases in beam size of even just small-sized LLM (1B-7B parameters) require an extensive GPU resource consumption, leading to up to 80% of recurring crashes due to memory overloads in LLM-based APR. Seemingly simple solutions to reduce memory consumption are (1) to quantize LLM models, i.e., converting the weights of a LLM from high-precision values to lower-precision ones. and (2) to make beam search sequential, i.e., forwarding each beam through the model sequentially and then concatenate them back into a single model output. However, we show that these approaches still do not work via both theoretical analysis and experiments.
To address this, we introduce FLAMES, a novel LLM-based APR technique that employs semantic-guided patch generation to enhance repair effectiveness and memory efficiency. Unlike conventional methods that rely on beam search, FLAMES utilizes greedy decoding to enhance memory efficiency while steering the search to more potentially good repair candidates via a semantic-guided best-first search algorithm. At each decoding step, FLAMES uses semantic feedback from test validation such as the number of passing and failing test cases to select the most promising token to explore further. Our empirical evaluation on the Defects4J and HumanEval-Java datasets shows that FLAMES not only substantially reduces memory consumption by up to 83% compared to conventional LLM-based APR, but also accelerates the repair process. Remarkably, FLAMES successfully generated 133 and 103 correct fixes for 333 and 163 bugs in the Defects4J and HumanEval-Java datasets, respectively. This suggests that FLAMES is not only more efficient but also outperforms state-of-the-art techniques, fixing at least 10 and 11 more bugs than SOTA baselines in the Defects4J and HumanEval-Java datasets, respectively.
Program Repair, Efficiency, Large Language Models
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software testing and debuggingccs: Computing methodologies Natural language processing
@@ 1. Introduction
Fixing software bugs is a complex and time-consuming task for developers (Winter et al., 2023). Automated Program Repair (APR)(Goues et al., 2019; Gao et al., 2022) has emerged as a promising solution to alleviate this burden by automatically repairing bugs. Over the past two decades, APR has seen significant advancements (Kim et al., 2013; Mechtaev et al., 2016; Le et al., 2016, 2017; Liu et al., 2019; Ghanbari et al., 2019; Chen et al., 2021a; Koyuncu et al., 2020; Le-Cong et al., 2021; Zhu et al., 2021; Xia and Zhang, 2022; Ye et al., 2022b), with practical applications in the software industry (Winter et al., 2022; Kirbas et al., 2021; Marginean et al., 2019; Bader et al., 2019). Recently, the rise of Large Language Models (LLMs), pre-trained on vast datasets, has further advanced APR, opening a new era of LLM-based APR (Silva et al., 2023; Xia et al., 2023b; Jiang et al., 2023a). LLM-based techniques typically adapt models for APR either through fine-tuning on APR-specific datasets (Silva et al., 2023; Jiang et al., 2023a; Huang et al., 2023; Paul et al., 2023; Zhang et al., 2022; Zirak and Hemmati, 2024) or by using prompting with commercial models (Bouzenia et al., 2024; Yin et al., 2024). In this study, we focus on fine-tuning approaches, as prompting often relies on closed-source and commercial models like ChatGPT, which lack transparency, hindering reproducibility in open science and raising data privacy concerns.
@@
Fine-tuning-based approaches typically refine the weights of Code LLMs, such as CodeLLama (Rozière et al., 2023) and Incoder (Fried et al., 2022) by fine tuning these models on APR datasets. During inference, they generate token sequences to construct candidate patches, prioritizing sequences with high probability scores from the fine-tuned model. Traditionally, beam search, a widely-used search algorithm from natural language processing (Sutskever et al., 2014), is used to optimize this process by maintaining only the top-$n$ nodes at each decoding step, where $n$ is the beam size. A larger beam size provides a more precise approximation to exact decoding, i.e., thoroughly exploring all possible sequences, thereby enhancing the quality of the outputs. This hyper-parameter is even more important in the context of APR, which covers a broader range of possible identifiers and a larger search space than natural language (Lutellier et al., 2020; Chen et al., 2021a; Jiang et al., 2021). It has been shown that a substantial beam size is essential for generating adequate candidate patches for optimal results (Jiang et al., 2021; Lutellier et al., 2020; Zhong et al., 2024), and a bigger beam size leads to higher performance of APR techniques that are based on small-sized deep learning models (less than 300M parameters)  (Tufano et al., 2019; Ye et al., 2022a; Zhong et al., 2024).
@@
But, is it true that increasing the beam size is all we need to improve the performance of APR? We revisit the impact of beam size on the effectiveness of LLM-based APR techniques on larger models (1B-7B parameters). Our findings reveal that increasing the beam size boosts the effectiveness of these techniques, but once the beam size is reached a threshold the performance drops significantly due to memory overloads that cause recurring crashes even on state-of-the-art hardware (NVIDIA A100 GPU, equipped with 80 GB of VRAM). For instances, state-of-the art LLM-based APR techniques such as InCoder (Jiang et al., 2023a), CodeGen (Jiang et al., 2023a), and RepairLLama (Silva et al., 2023) produced between 21% and 46% more plausible patches when the beam size was increased from 10 to 25. Ideally, we expect that further increasing the beam size would lead to better performance of the models for APR. However, when beam size is set higher at 50, 100, and 200, our experiments reveal that performance of these models significantly drops.
@@
Our experiments uncover that the above phenomenon is due to the extensive GPU resource consumption, particularly on Video Random Access Memory (VRAM), caused by larger beam sizes. Often, upgrading VRAM to accommodate larger beam sizes requires GPU upgrades, which are expensive, and thus this creates a costly trade-off between memory efficiency and the performance of LLM-based APRs. For instance, achieving a 46% performance improvement with InCoder-1B required 58% more average and 111% more peak VRAM usage on our dataset. Consequently, further performance gains through increasing beam size demand substantial VRAM, often leading to out-of-memory (OOM) crashes even on cutting-edge hardware and thereby unduly reducing the effectiveness of LLM-based APR techniques. Our experiments using the NVIDIA A100 GPU, equipped with 80 GB of VRAM, showed that Incoder-6B crashed on nearly 80% of the evaluated bugs when using a beam size of 200. This resulted in a 60% reduction in performance compared to a beam size of 10. In summary, we conclude that simply increasing the beam size is not a solution as it comes at the cost of chasing for state-of-the-art hardware.
@@
A seemingly straightforward approach to this problem is to reduce the memory usage of LLM-based program repair through engineering efforts. This can be achieved using two common strategies: (1) quantizing LLM models by converting their weights from high-precision to lower-precision values, and (2) making beam search sequential, where each beam is processed individually through the model before being combined into a single output. However, our findings show that while these methods can reduce memory usage, their memory demands still escalate significantly as beam size increases, leading to substantially high out-of-memory (OOM) rates. Consequently, these engineering efforts alone cannot solve the problem. This realization motivates us to develop a patch generation algorithm that can effectively enhance the performance of LLM-based APR techniques without relying on increasing the beam size and compromising memory efficiency.
@@
In this paper, we introduce FLAMES, an efficient LLM-based Program Repair technique using semantic-guided patch generation. Unlike conventional methods that rely solely on a language model’s knowledge and beam search, our approach aim to leverage semantic feedback from test validations to guide LLMs in patch generation. The core idea of FLAMES is to combine LLM-based and search-based APR. This process begins with generating initial patches using a greedy decoding, i.e., beam search with beam size of 1, followed by iterative, semantic-guided searches to refine these solutions. To achieve this, we employ a best-first search algorithm, specifically PG-TD (Zhang et al., 2023a), along with semantic feedback from test validations to guide the patch generation process in LLMs. FLAMES provides three primary advantages in addressing the limitations associated with beam search. First, it significantly reduces VRAM consumption by using greedy decoding, enhancing memory efficiency. Second, the approach scales well as it does not increase VRAM usage while generating more candidate patches. Third, FLAMES facilitates seamless information exchange between the patch generation and validation phases, enabling the efficient and effective exploration of plausible candidate patches.
@@
We conducted an empirical evaluation of our proposed approach, FLAMES, on a dataset including 333 real-world bugs from Defects4J and 163 synthetic bugs from the HumanEval dataset. Our experiments demonstrate that FLAMES successfully repaired 133 and 103 bugs from the Defects4J and HumanEval datasets, respectively. We compared FLAMES against 15 leading APR techniques. Our results indicate that FLAMES substantially outperforms these baselines, achieving at least 10 and 11 more correct fixes in the Defects4J and HumanEval datasets, respectively. We also found that FLAMES’s semantic-guided patch generation substantially outperforms widely-used patch generation strategies in LLM-based APR, including beam search and multiple sampling. The improvement rates range from 43% to 1145% across five different Large Language Models (LLMs) with three different architectures. Additionally, our analysis of memory efficiency showed that our method could reduce VRAM consumption by 42% to 83%, decreasing peak VRAM requirements from over 80 GB to as low as 12.7 GB across various configurations and models. Additionally, FLAMES also significantly boosts the time efficiency of LLMs, often generating plausible patches faster than conventional LLMs with beam search.
In summary, we made the following contributions:
@@
  • We introduce a novel approach, namely FLAMES, which fuses LLM-based and search-based APR, utilizing semantics feedback from patch validation to efficiently discover plausible candidate patches.
  • We empirically evaluate the performance of FLAMES against 15 state-of-the-art baselines. FLAMES correctly fixes 133 bugs from the Defects4J dataset and 103 bugs from the HumanEval-Java dataset, surpassing the best baseline by 10 and 11 fixes, respectively. FLAMES also outperforms widely-used patch generation strategies by 53% - 93%. Moreover, FLAMES substantially reduces memory consumption by up to 83% and accelerates the repair process.
@@ 2. Motivation Study
Beam size is a critical hyper-parameter in beam search, the fundamental algorithm for patch generation in LLM-based APR. In this section, we evaluate the impact of beam size on the effectiveness and memory efficiency of LLM-based APR techniques using LLMs with 1B to 7B parameters. While prior studies (Jiang et al., 2021; Lutellier et al., 2020; Xia and Zhang, 2022; Wei et al., 2023) showed that a substantial beam size is needed for optimal performance for APR, our experiments reveal that increasing the beam size unduly hinders memory efficiency and subsequently degrades the performance of LLM-based APR even on state-of-the-art hardware. This motivates a new search algorithm, which we propose in Section 3, to improve the efficiency and effectiveness of LLM-based APR. We structure our empirical study with the following two research questions:
  • RQ1: How does beam size affect memory efficiency of LLM-based APR? We measure VRAM usage and the frequency of out-of-memory crashes in LLM-based APR techniques with different beam sizes to understand the impact of this parameter on memory efficiency. We selected beam sizes of 10, 25, 50, 100, and 200, as these are common in prior works (Ye et al., 2022a; Xia et al., 2023b; Xia and Zhang, 2022; Zhu et al., 2021). We did not go beyond beam size of 200 due to hardware constraints.
  • RQ2: How does beam size affect the effectiveness of LLM-based APR? We investigate the impact of beam size on the effectiveness of 5 LLM-based APR techniques by measuring the number of plausible patches. Similarly to RQ1, we also selected beam sizes of 10, 25, 50, 100, and 200.
For both research questions, we also examine the impact of memory reduction techniques, as described in Section 2.1.3, on our findings.
@@ 2.1. Experimental Design
To address these research questions, we experiment on Defects4J, a well-established benchmark. This dataset was collected by Just et al.(Just et al., 2014) from 17 open-source Java projects, comprising 835 real-world bugs. In this study, following previous works(Ye et al., 2022a; Le-Cong et al., 2024; Xia and Zhang, 2022; Liu et al., 2019; Chen et al., 2021a), we focus on single-hunk bugs, where the patch is limited to a single contiguous chunk of code at a single location. We select this kind of bug as the existing LLM-based APR, which are studied in our work, such as InCoder and CodeGen (Jiang et al., 2023a), are only fine-tuned for single-hunk bugs (refer to next sections for details).In total, we identified 333 single-hunk bugs using the script provided by the original authors of Defects4J.
2.1.2. Studied LLM-based Program Repair techniques
To select the target techniques for our empirical study, we conducted a literature review and identified suitable techniques based on the following criteria: (1) use beam search as the patch generation strategy, (2) employ LLM with at least one billion parameters, (3) provide accessible fine-tuned models for Automated Program Repair.
Based on these criteria, we identified five LLM-based APR techniques from recent works (Jiang et al., 2023a; Silva et al., 2023) for our study including CodeGen-2B (Nijkamp et al., 2022), CodeGen-6B (Nijkamp et al., 2022), InCoder-1B (Fried et al., 2022), InCoder-6B (Fried et al., 2022), and RepairLlama (Silva et al., 2023). The CodeGen and InCoder models are LLMs which has been fine-tuned for Program Repair by Jiang et al.(Jiang et al., 2023a) and shown the best performance among evaluated techniques. Meanwhile, RepairLlama, developed by Silva et al.(Silva et al., 2023), builds on the CodeLLama (Rozière et al., 2023) model, fine-tuned for Program Repair using the QLoRA (Dettmers et al., 2024) method with optimized input and output representations. Note that, as RepairLlama has already used QLoRA and 8-bit quantization, we do not consider their quantization version of this technique in our follow-up experiments. Detailed information of these techniques is provided in our replication package (Authors, 2023).

### v2

# Memory-Efficient Large Language Models for Program Repair with Semantic-Guided Patch Generation

@@ Abstract.
Fixing software bugs is crucial yet demands significant resources from developers. Automated Program Repair (APR) is a promising solution to address this challenging task. The emergence of Large Language Models (LLMs) has opened a new era of LLM-based APR, substantially advancing the APR field further. LLM-based APR methods face significant challenges regarding memory inefficiency, hindering their scalability and effectiveness. This is largely due to the beam search utilized in the patch generation phase of LLM-based APR, which requires large beam sizes to search for more potentially good repair candidates.
In this paper, we first show that increases in beam size, even for small-sized LLMs (1B-7B params), require extensive GPU usage, leading to up to 80% of recurring crashes due to memory overloads in LLM-based APR. Seemingly simple solutions to reduce memory consumption are (1) to quantize LLM models, i.e., converting the weights of an LLM from high-precision values to lower-precision ones, and (2) to make beam search sequential, i.e., forwarding each beam through the model sequentially and then concatenating them back into a single output. However, we show that these approaches still do not work via both theoretical analysis and experiments.
To address this, we introduce FLAMES, a novel LLM-based APR technique that employs semantic-guided patch generation to enhance repair effectiveness and memory efficiency. Unlike conventional methods that rely on beam search, FLAMES utilizes greedy decoding to enhance memory efficiency while steering the search towards more potentially good repair candidates via a semantic-guided best-first search algorithm. At each decoding step, FLAMES uses semantic feedback from test validation, such as the number of passing and failing test cases, to select the most promising token to explore further. Our empirical evaluation on Defects4J shows that FLAMES substantially reduces memory consumption by up to 83% compared to LLM-based APR without compromising time efficiency. Moreover, FLAMES correctly fixes 133 bugs on Defects4J, fixing 10 bugs more than the best baseline. Additionally, these improvements also generalize to the HumanEval-Java and TransformedD4J datasets, where FLAMES generates 12% and 36.5% more correct patches, respectively, than the best baseline.
Program Repair, Memory Efficiency, Large Language Models
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 48th IEEE/ACM International Conference on Software Engineering; 12 - 18 April 2026; Rio De Janeiro, Brazilccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Fixing software bugs is a complex and time-consuming task for developers (Winter et al., 2023). Automated Program Repair (APR) (Goues et al., 2019; Gao et al., 2022) has emerged as a promising solution to alleviate this burden by automatically repairing bugs. Over the past two decades, APR has seen significant advancements (Le et al., 2016, 2017; Liu et al., 2019; Ghanbari et al., 2019; Chen et al., 2021a; Le-Cong et al., 2021; Zhu et al., 2021; Xia and Zhang, 2022; Ye et al., 2022b), with practical applications in the software industry (Winter et al., 2022; Kirbas et al., 2021; Marginean et al., 2019; Bader et al., 2019).
Recently, the surge of Large Language Models (LLMs), pre-trained on vast datasets, has further advanced APR, opening a new era of LLM-based APR (Silva et al., 2023; Xia et al., 2023b; Jiang et al., 2023a). LLM-based APR typically adapts models for APR either through fine-tuning on APR-specific datasets (Silva et al., 2023; Jiang et al., 2023a; Huang et al., 2023; Paul et al., 2023; Zhang et al., 2022; Zirak and Hemmati, 2024) or by using prompting with commercial models (Bouzenia et al., 2024; Yin et al., 2024). In this study, we focus on fine-tuning approaches, as prompting often relies on closed-source and commercial models like ChatGPT, which lack transparency, hindering reproducibility in open science and raising data privacy concerns.
@@
Fine-tuning-based approaches typically refine the weights of Code LLMs by fine-tuning these models on APR datasets. During inference, they generate token sequences to construct candidate patches, prioritizing sequences with high probability scores from the fine-tuned model. Traditionally, beam search, a widely used search algorithm from natural language processing (Sutskever et al., 2014), is used to optimize this process by maintaining only the top-$n$ nodes at each decoding step, where $n$ is the beam size. A larger beam size provides a more precise approximation to exact decoding, i.e., thoroughly exploring all possible sequences, thereby enhancing the quality of the outputs. This hyperparameter is even more important in the context of APR, which covers a broader range of possible identifiers and a larger search space than natural language (Lutellier et al., 2020; Chen et al., 2021a; Jiang et al., 2021). It has been shown that a substantial beam size is essential for generating adequate candidate patches for optimal results (Jiang et al., 2021; Lutellier et al., 2020; Zhong et al., 2024), and a larger beam size leads to higher performance of APR techniques that are based on small-sized deep learning models (less than 300M parameters)  (Tufano et al., 2019; Ye et al., 2022a; Zhong et al., 2024).
@@
But, is it true that increasing the beam size is all we need to improve the performance of APR? We revisit the impact of beam size on the effectiveness of LLM-based APR techniques on larger models (1B-7B parameters). Our findings reveal that increasing the beam size boosts the effectiveness of these techniques, but once the beam size reaches a threshold, the performance drops significantly due to memory overloads that cause recurring crashes even on state-of-the-art hardware (NVIDIA A100 GPU, equipped with 80 GB of VRAM). For instance, state-of-the-art LLM-based APR techniques generated between 21% and 46% more plausible patches when the beam size was increased from 10 to 25. Ideally, we expect that further increasing the beam size would lead to better performance of the models for APR. However, when the beam size is set higher at 50, 100, and 200, our experiments reveal that the performance of these models drops significantly.
@@
Our experiments uncover that the above phenomenon is due to extensive GPU resource consumption, particularly on Video Random Access Memory (VRAM), caused by larger beam sizes. Often, upgrading VRAM to accommodate larger beam sizes requires GPU upgrades, which are expensive; thus, this creates a costly trade-off between memory efficiency and the performance of LLM-based APRs. For instance, achieving a 46% performance improvement with InCoder-1B required 58% more average and 111% more peak VRAM usage on our dataset. Consequently, further performance gains through increasing beam size demand substantial VRAM, often leading to out-of-memory (OOM) crashes even on cutting-edge hardware and thereby unduly reducing the effectiveness of LLM-based APR techniques. Our experiments using the NVIDIA A100 GPU, equipped with 80 GB of VRAM, showed that InCoder-6B crashed on nearly 80% of the evaluated bugs when using a beam size of 200. This resulted in a 60% reduction in performance compared to a beam size of 10. In summary, we conclude that simply increasing the beam size is not a solution, as it comes at the cost of chasing after state-of-the-art hardware.
@@
A seemingly straightforward approach to this problem is to reduce the memory usage of LLM-based APR through engineering efforts. This can be achieved using two common strategies: (1) quantizing LLM models by converting their weights from high-precision to lower-precision values, and (2) making beam search sequential, where each beam is processed individually through the model before being combined into a single output. However, our findings show that while these methods can reduce memory usage, their memory demands still escalate significantly as beam size increases, leading to substantially high out-of-memory (OOM) rates. Consequently, these engineering efforts alone cannot solve the problem. This realization motivates us to develop a patch generation algorithm that can effectively enhance the performance of LLM-based APR techniques without relying on increasing the beam size and compromising memory efficiency.
@@
In this paper, we introduce FLAMES, a memory-efficient LLM-based APR technique using semantic-guided patch generation. Unlike conventional methods that rely solely on a language model’s knowledge and beam search, our approach aims to leverage semantic feedback from test validations to guide LLMs in patch generation. The core idea of FLAMES is to combine LLM-based and search-based APR. This process begins with generating initial patches using greedy decoding, i.e., beam search with a beam size of 1, followed by iterative, semantic-guided searches to refine these solutions. To achieve this, we employ a best-first search algorithm, specifically PG-TD (Zhang et al., 2023a), along with semantic feedback from test validations to guide the patch generation of LLMs. This approach offers three key advantages: (1) reduced VRAM consumption through greedy decoding, improving memory efficiency; (2) scalability without increasing VRAM usage while generating more candidate patches; and (3) seamless information exchange between patch generation and validation, enabling efficient exploration of plausible patches.
@@
We conducted an empirical evaluation of our proposed approach, FLAMES, using a dataset comprising 333 bugs from Defects4J (Just et al., 2014), 163 bugs from HumanEval-Java (Jiang et al., 2021), and 1,098 bugs from TransformedD4J (Le-Cong et al., 2025). We compared FLAMES against 15 state-of-the-art APR techniques and Qwen2.5-Coder-32B, the leading open-source LLM for code in well-known leaderboards (Liu et al., 2023; Zhuo et al., 2024). Our experimental results demonstrate that FLAMES can correctly fix 8%, 12%, and 36.5% more bugs than the best baseline on Defects4J, HumanEval-Java, and TransformedD4J, respectively. We also found that FLAMES’s semantic-guided patch generation substantially outperforms widely used patch generation strategies in LLM-based APR, including beam search and multiple sampling, with at least 12% and 23% improvements in the number of correctly fixed bugs. Moreover, our analysis of memory efficiency showed that our method could reduce VRAM consumption by 42% to 83%, decreasing peak VRAM requirements from over 80 GB to as low as 12.7 GB across various configurations and models. Despite the improvement in memory efficiency, FLAMES does not compromise time efficiency, even generating plausible patches faster than conventional LLMs using beam search in many cases.
In summary, we have made the following contributions:
@@
  • We introduce a novel approach, namely FLAMES, which fuses LLM-based and search-based APR, utilizing feedbacks from patch validation to efficiently discover plausible candidate patches.
  • We empirically evaluate the performance of FLAMES against 15 state-of-the-art baselines and the leading open-source LLM for code. FLAMES correctly fixes 133 bugs in Defects4J, 103 bugs in HumanEval-Java and 456 bugs in TransformedD4J, surpassing the best baseline by 8%, 12% and 36.5%, respectively. FLAMES also outperforms widely-used patch generation strategies by at least 12%. Moreover, FLAMES substantially reduces memory consumption by up to 83% without compromising time efficiency.
@@ 2. Motivation Study
Beam size is a critical hyperparameter in beam search, the fundamental algorithm for patch generation in LLM-based APR. In this section, we evaluate the impact of beam size on the effectiveness and memory efficiency of LLM-based APR techniques using LLMs. Although prior studies (Jiang et al., 2021; Lutellier et al., 2020; Xia and Zhang, 2022; Wei et al., 2023) showed that a substantial beam size is needed for optimal APR performance, our experiments reveal that increasing the beam size unduly hinders memory efficiency and subsequently degrades the performance of LLM-based APR, even on state-of-the-art hardware.
Models InCoder-1B InCoder-6B CodeGen-2B CodeGen-6B RepairLLama
Base Model InCoder (Fried et al., 2022) CodeGen-NL (Nijkamp et al., 2022) CodeLlama-7B (Rozière et al., 2023)
Pre-training Dataset InCoder (Fried et al., 2022) ThePile (Gao et al., 2020) CodeLLama (Rozière et al., 2023)
- Cut-off Date 12.2021 12.2020 08.2023
Fine-tuning Dataset Recoder (Zhu et al., 2021) MegaDiff (Monperrus et al., 2021)
- Cut-off Date 03.2018 09.2015
Fine-tuning Technique Full-parameter Fine-tuning QLORA (Dettmers et al., 2024)
Model Size 1B 6B 2B 6B 7B
Table 1. Detailed Information of LLM-based Program Repair techniques used in this study.
RQ1: How does beam size affect memory efficiency of the LLM-based APR? We measure the use of VRAM and the frequency of out-of-memory crashes in the LLM-based APR techniques with different beam sizes to understand the impact of this parameter on memory efficiency. We selected beam sizes of 10, 25, 50, 100, and 200, as these are common in prior works (Ye et al., 2022a; Xia et al., 2023b; Xia and Zhang, 2022; Zhu et al., 2021). We do not go beyond a beam size of 200 due to hardware constraints.
RQ2: How does beam size affect the effectiveness of LLM-based APR? We investigate the impact of beam size on the effectiveness of five LLM-based APR techniques by measuring the number of plausible patches. Similarly to RQ1, we also selected beam sizes of 10, 25, 50, 100, and 200.
@@ 2.1. Experimental Design
To address these research questions, we conducted experiments with Defects4J, a benchmark of 835 real-world bugs from 17 open-source Java projects (Just et al., 2014). Following prior studies (Ye et al., 2022a; Le-Cong et al., 2025; Xia and Zhang, 2022; Liu et al., 2019; Chen et al., 2021a), we focused on 333 single-hunk bugs, where patches modify a single contiguous code chunk. This selection aligns with LLM-based APR techniques like InCoder and CodeGen, which are fine-tune for single-hunk bug fixes (Jiang et al., 2023a).
2.1.2. Studied LLM-based APR techniques
To select the target techniques for our empirical study, we conducted a literature review and identified suitable techniques based on the following criteria: (1) use beam search as their patch generation strategy, (2) employ LLM with at least one billion parameters, and (3) provide accessible fine-tuned models for Automated Program Repair. Applying these criteria, we identified five LLM-based APR techniques from recent works (Jiang et al., 2023a; Silva et al., 2023): CodeGen (2B, 6B) and InCoder (1B, 6B), fine-tuned using full-parameter fine-tuning by Jiang et al.(Jiang et al., 2023a); and RepairLlama (7B), fine-tuned with efficient-parameter fine-tuning, i.e., QLORA, by Silva et al. (Silva et al., 2023). Detailed information about these LLMs are provided in Table 1. Since RepairLlama has utilized quantization, we excluded its quantized versions from follow-up experiments.